OpenStudy (anonymous):

what is the equation of the line that has a slope of and y-intercept of 2/3?

OpenStudy (anonymous):

y=mx+c m=slope,c=y intercept

OpenStudy (anonymous):

slope of ?????

OpenStudy (anonymous):

Slope of 0

OpenStudy (anonymous):

The y-intercept is 2/3?

OpenStudy (anonymous):

yes the y- intercept is 2/3

OpenStudy (anonymous):

Anyways, Slope-intercept form:\[y = mx + b\] where m is slope b is y-intercept \[y = (0)x + (2/3)\] \[y = (2/3)\]